OpenStudy (anonymous):

*I promise to medal and fan.* I'm horrible at trying to figure this out. Please help me! A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 7 experience points. On level 3, a player must earn 15 experience points. On level 4, a player must earn 25 experience points. Which recursive equation represents the pattern?

OpenStudy (anonymous):

A. an = an – 1 + n B. an = an – 1 + 3n + 1 C. an = an – 1 + 5n D. an = an – 1 + 2(n + 1)

OpenStudy (anonymous):

wuts an? a number?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

oh its actually supposed to look like this:

OpenStudy (anonymous):

like wat

OpenStudy (anonymous):

sorry taking a screenshot and i can't find it just a sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

haha, i have no idea how to do this.

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

hopefully he can help

OpenStudy (anonymous):

sorry :P

OpenStudy (anonymous):

its ok I appreciate you trying :)